When a negative integer is subtracted from another negative integer, the sign of the result (A) is always negative (B) is always positive (C) is never negative (D) depends on the numerical value of the integers

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

D) depends on the numerical value of the integers

For example: (i) - 4–(-3) = -4 + 3

= – 1
(ii)–6–(-9) = - 6 + 9

=3

Thanks